OpenStudy (coconut):

1. What is the 9th term of the geometric sequence 3, 12, 48, ...?

OpenStudy (anonymous):

ar^n r= 4 a=3 3(4)^(4)=768

OpenStudy (anonymous):

ar^n r= 4 a=3 3(4)^(9)=786432

OpenStudy (anonymous):

T9 = 3(4)^(9-1) = 196608

OpenStudy (coconut):

which is it :)?

OpenStudy (anonymous):

\[\LARGE a_n=a_1\cdot r^{n-1}\] \[\LARGE q=\frac{a_n}{a_{n-1}}=\frac{a_2}{a_1}=\frac{12}{3}=4\] so q=4 \[\LARGE a_9=3\cdot 4^{9-1}\] \[\LARGE a_9=3\cdot 4^8 \] \[\LARGE a_9=3\cdot 256\cdot 256 \] \[\LARGE a_9=3\cdot 65 536 \] \[\LARGE a_9=196608 \]

OpenStudy (anonymous):

sorry q=r